Exercise 7-16 Working with a Segmented Income Statement; Break-Even Analysis (LO7-4, LO7-5) [The following information applieRey IA кеч нь Rey 1 Compute the break-even point for the Chicago office and for the Minneapolis office. (Round CM ratio to
Answer #1

1A) Break even point = Fixed cost/CM ratio

= (126,000+63,000)/50%

= 378,000

1B)

Chicago = 78,000/70% = 111,429

Minneapolis = 48,000/40% = 120,000

Comment if you face any issues

​​​​​